# Epoch-based Optimistic Concurrency Control in Geo-replicated Databases

### v1

@@ 7. Conclusion
In this paper, we present Minerva, a novel replication solution for transactional geo-replicated distributed databases. Minerva leverages the benefits of Optimistic Concurrency Control (OCC) execution for scalable parallelism across nodes, combined with a novel, highly concurrent, and fault-tolerant replication protocol. Deterministic re-execution is used to ensure that transaction aborts are minimized without requiring additional coordination among replicas. Our experimental results show that Minerva outperforms state-of-the-art solutions, particularly in high-latency and high-replication-factor scenarios.

### v2

@@ 7. Conclusions
In this paper, we present Minerva, a novel distributed concurrency control protocol for transactional geo-replicated distributed databases. Minerva leverages the benefits of optimistic execution for scalable parallelism across nodes, combined with a novel, highly concurrent, and fault-tolerant replication protocol. Deterministic re-execution is used to ensure that transaction aborts are minimized without requiring additional coordination among replicas. Our experimental results show that Minerva outperforms state-of-the-art solutions, particularly in high-latency and high-replication-factor scenarios.
Acknowledgements.
This work was in part supported by NSERC and ORF. Generative AI tools such as ChatGPT were used to assist in the writing and editing of this paper.